This is a differential equations course, please answer all the questions.

Find the general solution.

2 11. :(t) -3) r(t) 12. z(t) = -1) =(4) 13. z(t) = -3) =(0) 14. z(t) = (-; -1)=(0) 2) -2 r(t) 5 10 15. z(t) = 2 1 3 2 16
Answer #1

con [8] [:)] dx, = 28,5x2 - at exp AP = 1, -222 - differentiate on wort t atp 2dx dt - 5 dx2 dt 11 2dxi dt (2x2-1x)st 27P ixplet D=m m²+1=0 m=ti 12,(t) = C, cost + C₂ sint differentiating q r 1 writt dt dt - 2dX2 dt X2 .(2x, -5%)-2022 dt2 dt 2 (de 1let D=m m²+1=0 m=ti X(t) = Cy cost + Cg sont XCH= G cost & + C₂ sint 424) cost & la Sint Xit) cost +C₂ Snit